# Fundamental Examples of non-compact Reeb spaces of smooth functions

### v1

@@ Abstract.
Reeb spaces of (continuous) real-valued functions on (nice) topological spaces are the spaces whose underlying sets consist of all connected components (contours) of their level sets and seen naturally as quotient spaces of the spaces. They are ”$1$-dimensional” spaces in various nice cases. They are graphs or graphs with ends for smooth function cases with nice singularities and behaviors. These tools or objects, have been fundamental and important in theory of Morse functions and more general smooth functions and applications to geometry, since the 20th century.
We present new systematic methods to obtain Reeb spaces homeomorphic to infinite graphs (with ends) for functions on non-compact manifolds with no boundary. This note is regarded as a note on cases previously obtained by the author from a kind of general viewpoints.
@@
Smooth, real analytic, or real algebraic (real polynomial) functions and maps. Reeb spaces. Graphs. Reeb graphs
@@
The Reeb space $R_{c}$ of a (continuous) real-valued function $c:X\rightarrow\mathbb{R}$ on a (nice) topological space $X$ is regarded to be the space of all connected components (contours) of all preimages (level sets) $c^{-1}(q)$. In the case of differentiable functions, we add ”regular” (resp. ”critical”) before ”contour” and ”level set” if it contains no critical point (resp. some critical points) of the function.
We introduce the notions mathematically and rigorously. We can induce the equivalence relation ${\sim}_{c}$ on $X$ by the rule: the relation $x_{1}{\sim}_{c}x_{2}$ holds if and only if $x_{1}$ and $x_{2}$ are in a same contour of a same level set $c^{-1}(y)$. The quotient map $q_{c}:X\rightarrow R_{c}:=X/{{\sim}_{c}}$ and the unique continuous function $\bar{c}:R_{c}\rightarrow\mathbb{R}$ with $c=\bar{c}\circ q_{c}$ is also defined.
@@
Hereafter, a graph means a $1$-dimensional connected CW complex such that the closure of each $1$-cell (edge) is homeomorphic to the $1$-dimensional disk $D^{1}:=\{x\in\mathbb{R}\mid x^{2}\leq 1\}$: 0-cell is a vertex. As a wider class, a graph with ends means a $1$-dimensional connected CW complex such that the closure of each $1$-cell (edge) is homeomorphic to the $1$-dimensional disk $D^{1}:=\{x\in\mathbb{R}\mid x^{2}\leq 1\}$ or a ray $\{x\geq 0\mid x\in\mathbb{R}\}$: 0-cell is a vertex as above.
In the case of the smooth manifold $X$ with no boundary, $R_{c}$ is, in a certain nice case, a graph (with ends) whose vertex set consists of all critical contours of $c$. This is the so-called Reeb graph of $c$ and in most studies including presented studies above, this definition is adopted. We can see the Reeb graph as a digraph (with ends) by giving orientation on each edge by the following rule.
  • If an edge $e$ departs from a vertex $v_{e,1}$, then it is an oriented edge departing from $v_{e,1}$.
  • If an edge $e$ enters a vertex $v_{e,2}$, then it is an oriented edge entering $v_{e,2}$.
@@
Can we construct a (nice) smooth function whose Reeb space has the structure of the Reeb graph (with ends) isomorphic to a given graph (resp. with ends)?
Hereafter, let ${\mathbb{R}}^{k}$ denote the $k$-dimensional Euclidean (real affine) space with $\mathbb{R}:={\mathbb{R}}^{1}$. Let ${\pi}_{k_{1},k_{2}}:{\mathbb{R}}^{k_{1}}\rightarrow{\mathbb{R}}^{k_{2}}$ denote the canonical projection mapping $x=(x_{1},x_{2})\in{\mathbb{R}}^{k_{1}-k_{2}}\times{\mathbb{R}}^{k_{2}}$ to $x_{1}$ with $k_{1}>k_{2}\geq 1$
@@
This is a kind of fundamental, natural and incredibly new problems. Sharko has launched in this 2006 ([17]). This is a study on smooth functions on closed surfaces whose critical points $p$ are of the form $c(z)={\Re}z^{l}+t_{p}$ or critical points of Morse functions and whose Reeb graphs are given graphs of a certain class. Here $\mathbb{C}$ is the space of all complex numbers, $l>1$ is an integer, $t_{p}\in\mathbb{R}$, and ${\Re}s$ denotes the real part of $s\in\mathbb{C}$. This is extended to arbitrary finite graphs in [12] and functions whose critical points may not be isolated are studied. In [13], a certain class of finite graphs is important and Morse functions on closed manifolds with prescribed Reeb graphs of the class and regular contours which are spheres are reconstructed. The author has contributed to this, in, [3, 5], for example. There the author respects not only graphs, but also shapes of level sets. In addition, in [4], the author considers smooth functions on non-compact manifolds and so-called non-proper functions, explicitly, first. As another important pioneering study by the author, [6] is on real algebraic construction with prescribed Reeb graphs. The author is contributing to real algebraic construction, mainly, recently, where we ignore related knowledge, and omit precise exposition on other related papers and preprints, except [7, 8]. Later, as a kind of our main ingredients, related arguments appear in our main ingredients and we discuss the arguments in self-contained ways.
@@
Recently the author has started to consider infinite graphs (with ends) as Reeb spaces and obtained several explicit cases. They are also constructed as, fully or densely, real algebraic or real analytic and based on the paper [5] and the preprints [7, 8]. Related preprints are [9, 10]. These preprints are important in our present paper, where we ignore related non-trivial knowledge and discuss related arguments in self-contained ways. The present paper is also a kind of our new remark on examples of [9, 10], especially [9, Theorem 1]. There real algebraic or real analytic functions or maps are presented and an answer to Problem 1 are presented as the restrictions of the canonical projections ${\pi}_{k_{1},k_{2}}$ to the zero sets of real analytic or algebraic maps on the real affine spaces. We respect these examples and present an answer from a kind of essentially generalized viewpoints.
In the next section, we present our main result and prove the new theorem (Theorem 3). This is for a generalization of the case presented in [9, Theorem 2] and in Example 1, this is explained again.
@@
For a differentiable map between differentiable manifolds, a point in the manifold of the domain is a singular point of it where the rank of the differential of the map is smaller than both the dimension of the manifold of the domain and that of the manifold of the target. The singular set of the map is defined as the set of all singular points of the map. We use ”critical” in the case where the map is the real-valued map.
@@
    The function $c_{i}:\mathbb{R}\rightarrow\mathbb{R}$ ($i=1,2$) is a smooth function whose critical set is a discrete set and at least one function of $c_{1}$ and $c_{2}$ has at least one critical point. Suppose also the following.
@@
      $t\in\mathbb{R}$ goes to -$\infty$, then the values $c_{1}(t)$ and $c_{2}(t)$ go to $b_{1}\in\mathbb{R}$.
@@
      $t\in\mathbb{R}$ goes to $\infty$, then the values $c_{1}(t)$ and $c_{2}(t)$ go to $b_{2}\in\mathbb{R}$.
@@
      For $i=1,2$, each connected component $K_{c_{1},c_{2},b_{i},j}$ of the set ${{\pi}_{2,1}}^{-1}(b_{i})\bigcap\overline{D_{c_{1},c_{2}}}$ has a small compact and connected neighborhood $N_{c_{1},c_{2},b_{i},j}$ consisting of contours of the the restriction ${\pi}_{2,1}{\mid}_{\overline{D_{c_{1},c_{2}}}}$ and the complementary set of $K_{c_{1},c_{2},b_{i},j}$ in the connected component $N_{c_{1},c_{2},b_{i},j}$ consists of finitely many connected sets. In addition, these neighborhoods $N_{c_{1},c_{2},b,j}$ can be chosen to be mutually disjoint.
@@
We prove the statement (2). First, the critical sets of the functions $c_{1}$ and $c_{2}$ are discrete by the assumption. From the conditions (2a, 2b) and the structures of the maps and the manifolds, the restriction of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ to the preimage of $\mathbb{R}-\{b_{1},b_{2}\}$ is proper. The critical set of the restriction of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ to the preimage of each closed subset in $\mathbb{R}-\{b_{1},b_{2}\}$ is finite, and by virtue of [15, 16], its Reeb space is regarded to be homeomorphic to a $1$-dimensional manifold or a graph which is not homeomorphic to a manifold. In addition, if at least one critical point of the function exists, then the Reeb space of the restriction of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ to the preimage of $\mathbb{R}-\{b_{1},b_{2}\}$ is regarded to be the Reeb graph. For [15, 16], related to this argument, mainly, consult [15, Theorem 3.1] and [16, Theorems 2.1 and 2.8 and ”2 Reeb space and its graph structure”]. Related to this theory, see also [1]. We consider the condition (2c). According to this, around each point in ${\bar{{\pi}_{2,1}{\mid}_{\overline{D_{c_{1},c_{2}}}}}}^{-1}(\{b_{1},b_{2}\})$, the Reeb space is homeomorphic to a $1$-dimensional manifold or a finite graph which is not homeomorphic to a manifold and regarded to be the Reeb graph (if there exists a critical point of the function ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ there), locally. This also comes from the general theory above.
@@
Theorem 3 is our main result.
Theorem 3.
@@
We choose a point $p$ of $X_{m,c_{1},c_{2}}$ mapped to a point $p_{e}$ in some edge $e$ of the Reeb graph by the quotient map and a point $p_{D}:={\pi}_{2,1}(p)\in\overline{{D}_{m,c_{1},c_{2}}}$. Let $p_{D}:=(p_{D,1},p_{D,2})$. We consider the point $p_{D,c_{2}}:=(c_{2}(p_{D,2}),p_{D,2})$. The set $\{(t,p_{D,2})\mid p_{D,1}\leq t\leq c_{2}(p_{D,2})\}$ is in $\overline{{D}_{m,c_{1},c_{2}}}$. Due to the assumption (2), we can consider some path from $p_{D,c_{2}}$ to another point $p_{s,c_{2}}$ in $\overline{{D}_{m,c_{1},c_{2}}}-{D}_{m,c_{1},c_{2}}$ and have the path departing from $p_{D}$ obtained by joining the straight segment $\{(t,p_{D,2})\mid p_{D,1}\leq t\leq c_{2}(p_{D,2})\}$ and the path in $\overline{{D}_{m,c_{1},c_{2}}}-{D}_{m,c_{1},c_{2}}$ such that the restriction of ${\pi}_{2,1}$ there is injective and that the point $(p_{s,c_{2}},{(0)}_{j=1}^{m-1})\in{{\mathbb{R}}}^{2}\times{\mathbb{R}}^{m-1}={\mathbb{R}}^{m+1}$ is a critical point of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$. Related to the assumption (1), we can argue similarly. We can see that the closure of $e$ in the Reeb graph is homeomorphic to $D^{1}$. More precisely, from the argument above, we naturally have a union of finitely many closures of edges of the Reeb graph with the union containing $p_{e}$. ∎
@@
    The case $c_{1}(x):=\frac{1}{2(x^{2}+1)}{\sin}^{2}x$ and $c_{2}(x):=\frac{1}{(x^{2}+1)}$ is an explicit case of Theorem 3 with the critical set of $c_{1}$ being unbounded, $c_{2}$ being almost unimodal, and $(b_{1},b_{2})=(0,0)$. This is also [9, Theorem 1].

### v2

# Fundamental examples of Reeb spaces of smooth functions defined from two graphs of smooth functions with same asymptotic behaviors

@@ Abstract.
Reeb spaces of (continuous) real-valued functions on (nice) topological spaces are the spaces whose underlying sets consist of all connected components (contours) of their level sets and seen naturally as quotient spaces of the spaces. They are ”$1$-dimensional” spaces in various nice cases. They are graphs or graphs with ends for smooth function cases with nice singularities and behaviors. Reeb spaces have been fundamental and important in theory of Morse functions and more general smooth functions and applications to geometry, since the 20th century.
We present Reeb spaces homeomorphic to infinite graphs (with ends) for functions on non-compact manifolds with no boundary. This paper is a note on cases previously obtained by the author. More explicitly, we consider a natural smooth map onto the region surrounded by the graphs of two smooth real-valued functions in the plane and its composition with the canonical projection.
@@
Smooth, real analytic, or real algebraic (real polynomial) functions and maps. Reeb spaces. Graphs. Digraphs. Reeb graphs
@@
The Reeb space $R_{c}$ of a (continuous) real-valued function $c:X\rightarrow\mathbb{R}$ on a (nice) topological space $X$ is regarded to be the space of all connected components (contours) of all preimages (level sets) $c^{-1}(q)$. In the case of differentiable functions, we add ”regular” (resp. ”critical”) before ”contour” and ”level set” if it contains no critical point (resp. some critical points) of the function. We explain the notions mathematically and rigorously. We can induce the equivalence relation ${\sim}_{c}$ on $X$ by the following rule: the relation $x_{1}{\sim}_{c}x_{2}$ holds if and only if $x_{1}$ and $x_{2}$ are in a same contour of a same level set $c^{-1}(y)$. The quotient map $q_{c}:X\rightarrow R_{c}:=X/{{\sim}_{c}}$ and the unique continuous function $\bar{c}:R_{c}\rightarrow\mathbb{R}$ with $c=\bar{c}\circ q_{c}$ is also defined.
@@
Hereafter, let ${\mathbb{R}}^{k}$ denote the $k$-dimensional Euclidean (real affine) space with $\mathbb{R}:={\mathbb{R}}^{1}$. Let ${\pi}_{k_{1},k_{2}}:{\mathbb{R}}^{k_{1}}\rightarrow{\mathbb{R}}^{k_{2}}$ denote the canonical projection mapping $x=(x_{1},x_{2})\in{\mathbb{R}}^{k_{1}-k_{2}}\times{\mathbb{R}}^{k_{2}}$ to $x_{1}$ with $k_{1}>k_{2}\geq 1$. A graph means a $1$-dimensional connected CW complex such that the closure of each $1$-cell (edge) is homeomorphic to the $1$-dimensional disk $D^{1}:=\{x\in\mathbb{R}\mid x^{2}\leq 1\}$: 0-cell is a vertex. As a wider class, a graph with ends means a $1$-dimensional connected CW complex such that the closure of each $1$-cell (edge) is homeomorphic to the $1$-dimensional disk $D^{1}:=\{x\in\mathbb{R}\mid x^{2}\leq 1\}$ or a ray $\{x\geq 0\mid x\in\mathbb{R}\}$: 0-cell is a vertex as above.
In the case of the smooth manifold $X$ with no boundary, $R_{c}$ is, in a certain nice case, a graph (with ends) whose vertex set consists of points for all critical contours of $c$. This is the so-called Reeb graph of $c$ and in most studies including presented studies above, this definition is adopted. We can see the Reeb graph as a digraph (with ends) by giving orientation on each edge by the following rule: an edge $e$ departing from a vertex $v_{e}$ is an oriented edge departing from (resp. entering) $v_{e}$.
@@
Can we construct a (nice) smooth function whose Reeb space is regarded to be the Reeb graph isomorphic to a given graph (resp. with ends)?
@@
This is a kind of fundamental, natural and incredibly new problems. Sharko has launched in this 2006 ([17]). This is a study on smooth functions on closed surfaces whose critical points $p$ are of the form $c(z)={\Re}z^{l}+t_{p}$ or critical points of Morse functions and whose Reeb graphs are given graphs of a certain class. Here $\mathbb{C}$ is the space of all complex numbers, $l>1$ is an integer, $t_{p}\in\mathbb{R}$, and ${\Re}s$ denotes the real part of $s\in\mathbb{C}$. This is extended to arbitrary finite graphs in [12] and functions whose critical points may not be isolated are studied. In [13], a certain class of finite graphs is important and Morse functions on closed manifolds with prescribed Reeb graphs of the class and regular contours which are spheres are reconstructed. The author has contributed to this, in, [3, 5], for example. There the author respects not only graphs, but also shapes of level sets. In addition, in [4], the author considers smooth functions on non-compact manifolds and so-called non-proper functions, explicitly, first. As another important pioneering study by the author, [6] is on real algebraic construction with prescribed Reeb graphs. The author is contributing to real algebraic construction, mainly, recently, where we ignore related knowledge, and omit precise exposition on other related papers and preprints, except [7, 8]. Later, as a kind of our main ingredients, related arguments appear in our main ingredients and we discuss the arguments in self-contained ways (Theorem 1).
@@
Recently the author has started to consider infinite graphs (with ends) as Reeb spaces and obtained several explicit cases. They are also constructed as, fully or densely, real algebraic or real analytic and based on the paper [5] and the preprints [7, 8]. Related preprints are [9, 10]. These preprints are important in our present paper, where we ignore related non-trivial knowledge and discuss related arguments in self-contained ways. The present paper is also a kind of our new remark on examples of [9, 10], especially [9, Theorem 1]. There real algebraic or real analytic functions or maps are presented and answers to Problem 1 are presented as the restrictions of the canonical projections ${\pi}_{k_{1},k_{2}}$ to the zero sets of real analytic or algebraic maps on the real affine spaces. We respect these examples and present answers from a kind of essentially generalized viewpoints.
In the next section, we prove our new propositions and theorems there. This is for a generalization of the case presented in [9, Theorem 2] and in Example 1, this is explained again. Explicitly, we consider a natural smooth map onto the region surrounded by the graphs of two smooth real-valued functions $c_{1}$ and $c_{2}$ goes to a same value at each infinity in ${\mathbb{R}}^{2}$, obtained via Theorem 1, and its composition with the canonical projection ${\pi}_{2,1}$.
@@
For a differentiable map between differentiable manifolds, a point in the manifold of the domain is a singular point of it where the rank of the differential of the map is smaller than both the dimension of the manifold of the domain and that of the manifold of the target. The singular set $S(c)$ of the map $c:X\rightarrow Y$ is defined as the set of all singular points of the map. The singular value set of the map $c$ is the image $c(S(c))$ of the singular set of the map. We use ”critical” in the case where the map is the real-valued map ($Y:=\mathbb{R}$).
@@
    Suppose that the function $c_{i}:\mathbb{R}\rightarrow\mathbb{R}$ ($i=1,2$) is a smooth function the intersection of $\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\})$ and whose critical value set $c_{i}(S(c_{i}))$ is a discrete set and at least one function of $c_{1}$ and $c_{2}$ has at least one critical point. Suppose also the following.
@@
      As $t\in\mathbb{R}$ goes to -$\infty$, then both the values $c_{1}(t)$ and $c_{2}(t)$ converge to $q_{1}\in\mathbb{R}$.
@@
      As $t\in\mathbb{R}$ goes to $\infty$, then both the values $c_{1}(t)$ and $c_{2}(t)$ converge to $q_{2}\in\mathbb{R}$.
@@
      For $i=1,2$, each connected component $K_{c_{1},c_{2},q_{i},j}$ of the set ${{\pi}_{m+1,1}}^{-1}(q_{i})\bigcap X_{m,c_{1},c_{2}}$ has a small compact and connected neighborhood $N_{c_{1},c_{2},q_{i},j}$ consisting of contours of the the restriction ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ and the complementary set of $K_{c_{1},c_{2},q_{i},j}$ in the connected component $N_{c_{1},c_{2},q_{i},j}$ consists of finitely many connected sets and contains no critical contour of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$. In addition, these neighborhoods $N_{c_{1},c_{2},q_{i},j}$ can be chosen to be mutually disjoint.
@@
We prove the statement (2). First, the critical sets of the functions $c_{1}$ and $c_{2}$ are discrete by the assumption. From the conditions (2a, 2b) and the structures of the maps and the manifolds, the restriction of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ to the preimage of $\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\})$ is proper. The critical set of the restriction of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ to the preimage of each closed subset in $\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\})$ is finite, and by virtue of [15, 16], its Reeb space is regarded to be homeomorphic to a $1$-dimensional manifold or a graph which is not homeomorphic to a manifold. In addition, if at least one critical point of the function exists, then the Reeb space of the restriction of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ to the preimage of $\mathbb{R}-(\{q_{1}\}\bigcup\{q_{2}\})$ is regarded to be the Reeb graph. For [15, 16], related to this argument, mainly, consult [15, Theorem 3.1] and [16, Theorems 2.1 and 2.8 and ”2 Reeb space and its graph structure”]. Related to this theory, see also [1]. We consider the condition (2c). According to this, around each point in ${\bar{{\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}}}^{-1}(\{q_{1}\}\bigcup\{q_{2}\})$, the Reeb space is homeomorphic to a $1$-dimensional manifold or a finite graph which is not homeomorphic to a manifold and regarded to be the Reeb graph (if there exists a critical point of the function ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ there), locally. This also comes from the general theory above.
@@
Proposition 1 is seen as a part of our main result.
Proposition 1.
@@
We choose a point $p$ of $X_{m,c_{1},c_{2}}$ mapped to a point $p_{e}$ in some edge $e$ of the Reeb graph by the quotient map and a point $p_{D}:={\pi}_{2,1}(p)\in\overline{{D}_{m,c_{1},c_{2}}}$. Let $p_{D}:=(p_{D,1},p_{D,2})$. We consider the point $p_{D,c_{2}}:=(c_{2}(p_{D,2}),p_{D,2})$. The set $\{(t,p_{D,2})\mid p_{D,1}\leq t\leq c_{2}(p_{D,2})\}$ is in $\overline{{D}_{m,c_{1},c_{2}}}$. Due to the assumption (2), we can consider some path from $p_{D,c_{2}}$ to another point $p_{s,c_{2}}$ in $\overline{{D}_{m,c_{1},c_{2}}}-{D}_{m,c_{1},c_{2}}$ and have the path departing from $p_{D}$ obtained by joining the set $\{(t,p_{D,2})\mid p_{D,1}\leq t\leq c_{2}(p_{D,2})\}$, which is a one-point set or a segment, and the path in $\overline{{D}_{m,c_{1},c_{2}}}-{D}_{m,c_{1},c_{2}}$ in such a way that the restriction of ${\pi}_{2,1}$ to the resulting path is injective and that the point $(p_{s,c_{2}},{(0)}_{j=1}^{m-1})\in{{\mathbb{R}}}^{2}\times{\mathbb{R}}^{m-1}={\mathbb{R}}^{m+1}$ is a critical point of ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$. Related to the assumption (1), we can argue similarly. We can see that the closure of $e$ in the Reeb graph is homeomorphic to $D^{1}$. More precisely, from the argument above, we naturally have a union of finitely many closures of edges of the Reeb graph with the union containing $p_{e}$. ∎
@@
    The case $c_{1}(x):=\frac{1}{2(x^{2}+1)}{\sin}^{2}x$ and $c_{2}(x):=\frac{1}{x^{2}+1}$ is an explicit case of Proposition 1 with the critical set of $c_{1}$ being unbounded, $c_{2}$ being (almost) unimodal, and $(q_{1},q_{2})=(0,0)$. This is also [9, Theorem 1].